OpenStudy (anonymous):

Divide and check: (36t^6 − 27t^5 − 9t^2) ÷ (−9t^2)

OpenStudy (saifoo.khan):

-4t^4 + 3t^3 +1

OpenStudy (anonymous):

how would the steps look like

OpenStudy (saifoo.khan):

it's mental though.

OpenStudy (anonymous):

huh?

OpenStudy (saifoo.khan):

There is no particular working for it, i solved it mentally.

OpenStudy (saifoo.khan):

But if u want to show working, then \[\frac{36t^6}{-9t^2} -\frac{ 27t^5}{ − 9t^2}+\frac{-9t^2}{-9t^2}\]

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

Welcome.